A soccer ball is kicked at an angle of 22 degrees and a velocity of 30 m/s. What is the highest point the ball reaches?


1
Expert's answer
2020-11-23T10:27:33-0500

Let's write the equation of motion of the soccer ball in vertical direction:


"y=v_0t_{rise}sin\theta-\dfrac{1}{2}gt_{rise}^2,"


here, "v_0=30\ \dfrac{m}{s}" is the initial velocity of the soccer ball, "t_{rise}" is the time that the soccer ball takes to reach the maximum height (or the highest point), "\theta=22^{\circ}" is the launch angle, "y" is the vertical displacement of the soccer ball (or the height) and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the soccer ball takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."


Then, we can substitute "t_{rise}" into the equation for "y" and find the maximum height:



"y_{max}=v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=\dfrac{(30\ \dfrac{m}{s})^2\cdot sin^222^{\circ}}{2\cdot 9.8\ \dfrac{m}{s^2}}=6.44\ m."

Answer:

"y_{max}=6.44\ m."